# Compounding Vulnerability: Hub Removal Triggers Cascade Phase Transitions While Degrading Percolation Robustness in Scale-Free Networks

### 05 Mar 2026 v1

@@ V.3 Derivation of $z_{1}$ After Hub Removal
For $m=2$, $k_{\text{cut}}=6$: expansion factor =8/5=1.6=8/5=1.6, predicting the cascade window expands from $\varphi^{*}\approx 0.20$ to $\varphi^{*}\approx 0.32$ (mean-field) or ${\sim}0.26$ (exact binomial).
@@ V.4 Numerical Verification of $z_{1}$
The exact binomial computation (not the mean-field approximation) confirms the mechanism:
@@
At $\varphi=0.22$, the pre-removal network has $z_{1}=0.850<1$ (subcritical: cascades die) while the post-removal network has $z_{1}=1.195>1$ (supercritical: global cascades possible). This crossing explains the observed cascade explosion. The newly opened cascade window spans $\varphi\in(0.20,{\sim}0.26)$, precisely containing the observed phase transition.
@@ V.5 Percolation Threshold Increase
Proposition 3 (Hub removal increases $p_{c}$). For any network $G$ with degree heterogeneity $\kappa>2$, hub removal of fraction $\alpha>0$ increases the Molloy–Reed bond percolation threshold.
This follows from the Molloy–Reed criterion $p_{c}=1/(\kappa-1)$ [18]: hub removal selectively removes nodes contributing $k^{2}$ to $\langle k^{2}\rangle$, reducing $\kappa$ and therefore increasing $p_{c}$. The result is exact for uncorrelated configuration-model networks and holds directionally for BA networks with weak degree-degree correlations.
@@ VI.1 The Dual-Function Constraint
Our results reveal that within the intervention class of uniform node deletion on static, unweighted networks, hub nodes simultaneously provide percolation robustness (via their $k^{2}$ contribution to $\langle k^{2}\rangle$, keeping $p_{c}$ low) and cascade resistance (via their high activation barrier $k\cdot\varphi\gg 1$, blocking cascade propagation). Both functions are monotonically linked through the single structural parameter $k$. Within this intervention class, there is no way to modify one function without simultaneously modifying the other. Alternative intervention classes—degree-preserving rewiring, weighted edge adjustment, or adaptive threshold modification—may decouple these functions and are an important direction for future work.
@@ VI.2 Relation to Prior Work
Watts (2002) [7] described stable nodes conceptually; we operationalize this as the quantitative condition $k>1/\varphi$ and demonstrate its dynamical nature through controlled experiments. The hub vulnerability experiment (§III) provides, to our knowledge, the first direct experimental isolation of the stable-node mechanism from connectivity effects in the context of cascade dynamics.
@@
Gleeson (2008) [24] extended analytical cascade methods to correlated and modular networks in the static (unperturbed) case, providing the natural framework for extending our $z_{1}$ derivation to networks with degree-degree correlations—a direction we identify as important future work.
Morone and Makse (2015) [15] and Braunstein et al. (2016) [16] optimize network dismantling for percolation fragmentation. Our results complement this literature by identifying the cascade side-effects of percolation-optimal interventions: dismantling strategies that efficiently fragment a network may simultaneously expand the cascade vulnerability window [26].
VI.3 Applicability to Real-World Networks
Our results predict compounding vulnerability in networks with degree heterogeneity $\kappa>{\sim}10$. Many real-world networks exceed this threshold: yeast protein-protein interaction networks ($\kappa\approx 17$ [22]), human PPI networks ($\kappa\approx 28$), the World Wide Web ($\kappa\approx 55$ [23]), and citation networks ($\kappa>100$). Conversely, power grids ($\kappa\approx 2$) and AS-level Internet ($\kappa\approx 3.5$) fall below the threshold, consistent with our prediction of no cascade phase transition in homogeneous networks.
@@ VI.5 Limitations
(i) Results are demonstrated on synthetic networks; empirical validation on real-world networks with $\kappa>10$ is needed. (ii) The $S=0.5$ percolation threshold estimator, while intuitive, is nonstandard; we note that the qualitative result ($\Delta F_{p}\gg 0$) is robust to estimator choice, as the Molloy–Reed analytical estimates (Appendix A) confirm the same directional effect despite quantitative differences. (iii) Uniform cascade thresholds are a simplification; heterogeneous thresholds may shift $\varphi^{*}$ but preserve the qualitative mechanism. (iv) The configuration-model approximation in §V.3 ignores BA disassortativity and the correlation structure induced by targeted removal; a full treatment using degree-dependent edge retention probabilities would strengthen the analytical claims, though the numerical match between the approximate $z_{1}$ predictions and simulation results suggests the approximation captures the dominant mechanism. (v) Static networks are studied; adaptive rewiring could mitigate or amplify the effect. (vi) The Watts cascade model is one of several cascade mechanisms; generalization to load-redistribution [17], cascade-based attacks [20], and DebtRank [13] models is an important direction. (vii) The $\kappa>{\sim}10$ threshold is derived for BA networks; the precise threshold depends on both the degree distribution shape and the cascade threshold $\varphi$, and should be understood as a BA-specific onset condition rather than a distribution-independent constant.
@@ VII Conclusion
We have demonstrated that hub removal in Barabási–Albert scale-free networks produces compounding vulnerability: it raises the bond percolation threshold by $168\%$ ($p_{c}$: $0.34\to 0.90$) while simultaneously expanding the Watts cascade window, enabling global cascades at threshold values that were previously subcritical. The phase transition at $\varphi=0.22$0.250.25—where mean cascade size jumps from $1\%$ to ${\sim}20\%$ of the network—is the most striking consequence, and it is silent during normal operation: subcritical cascade behavior is observationally indistinguishable from safe operation until the intervention shifts the system across the phase boundary.
@@
AI Assistance Disclosure
This work was conducted as a human–AI collaboration. The author conceived the research question, designed the hub vulnerability experiment, identified the cascade firewall mechanism, directed the analysis, and made all scientific decisions including the reframing from “trade-off” to “compounding vulnerability.” AI tools assisted with simulation code generation, manuscript drafting, and verification checks. The manuscript was subjected to multiple adversarial LLM review passes, and issues found during review were addressed in revision. All simulation code and random seeds are provided for independent reproduction.
@@ Acknowledgements.
The author thanks the open-source Python community for NetworkX and NumPy.
@@ Appendix A Percolation Method Comparison
Table 6: Comparison of simulation-based and Molloy–Reed $p_{c}$ estimates.
Network $p_{c}$ (sim) $p_{c}$ (MR) Ratio
[…]
@@ Appendix B $z_{1}$ Derivation Details
Appendix C Simulation Code Availability
@@
A Lean 4 formalization of selected analytical claims (e.g., stable-node/firewall inequalities used in §V) is provided in proofs/ in the same repository.
Key scripts: simulate_production.py (finite-size scaling), phi_sweep_large.py ($\varphi$-sweep), hub_vulnerability_test.pyIII experiment).

### 07 Mar 2026 v2

@@ V.3 Derivation of $z_{1}$ After Hub Removal
For $m=2$, $k_{\text{cut}}=6$: expansion factor =8/5=1.6=8/5=1.6, predicting the cascade window expands from $\varphi^{*}\approx 0.20$ to $\varphi^{*}\approx 0.32$ under the mean-field approximation. Equation (4) is a mean-field prediction that overestimates the true cascade window boundary because it averages over the integer-valued vulnerability cutoff $\lfloor 1/\varphi\rfloor$.
@@ V.4 Numerical Verification of $z_{1}$
To assess the accuracy of the mean-field approximation, we compute $z_{1}$ exactly using the post-removal degree distribution $P^{\prime}(j)$ from Eq. (2) with binomial edge retention. These exact values, which we use for all quantitative claims throughout the paper, give a tighter cascade window boundary of $\varphi^{*}_{\mathrm{post}}\approx 0.26$, in good agreement with simulation:
@@
At $\varphi=0.22$, the pre-removal network has $z_{1}=0.850<1$ (subcritical: cascades die) while the post-removal network has $z_{1}=1.195>1$ (supercritical: global cascades possible). This crossing explains the observed cascade explosion. The newly opened cascade window spans $\varphi\in(0.20,{\sim}0.26)$, precisely matching the observed Regime II boundaries.
Sensitivity to $\rho$ estimation. The configuration-model approximation gives $\rho=0.375$, while direct measurement on BA networks yields $\rho_{\mathrm{emp}}=0.421\pm 0.011$ (12.3% higher), reflecting degree–degree correlations absent in the configuration model. Using $\rho_{\mathrm{emp}}$, the post-removal $z_{1}$ at $\varphi=0.22$ decreases from 1.195 to 1.140—still supercritical ($z_{1}>1$). The crossing $z_{1}^{\mathrm{pre}}<1<z_{1}^{\mathrm{post}}$ is preserved, confirming that the cascade phase transition prediction is robust to the $\rho$ underestimation inherent in the configuration-model approximation.
@@ V.5 Percolation Threshold Increase
Proposition 1 (Hub removal increases $p_{c}$ in uncorrelated networks). For any uncorrelated random network (configuration model) with degree heterogeneity $\kappa>2$, hub removal of fraction $\alpha>0$ increases the Molloy–Reed bond percolation threshold.
This follows from the Molloy–Reed criterion $p_{c}=1/(\kappa-1)$ [19]: hub removal selectively removes nodes contributing $k^{2}$ to $\langle k^{2}\rangle$, reducing $\kappa$ and therefore increasing $p_{c}$. The result is exact for uncorrelated configuration-model networks and holds directionally for BA networks, whose weak disassortativity produces only a small quantitative correction (§V.3).
@@ VI.1 The Dual-Function Constraint
Our results reveal that within the intervention class of uniform node deletion on static, unweighted networks, hub nodes simultaneously provide percolation robustness (via their $k^{2}$ contribution to $\langle k^{2}\rangle$, keeping $p_{c}$ low) and cascade resistance (via their high activation barrier $k\cdot\varphi\gg 1$, blocking cascade propagation). Both functions are monotonically linked through the single structural parameter $k$. Across all tested networks and parameter combinations ($\varphi\in[0.15,0.28]$, $\alpha=0.10$, BA/ER/WS topologies, and four real-world networks), we observe no Pareto improvement: uniform hub deletion worsens both metrics simultaneously in every case examined. We note that this is an empirical observation, not a proven impossibility result; whether alternative intervention classes—degree-preserving rewiring, weighted edge adjustment, or adaptive threshold modification—can decouple these functions is an important open question for future work.
@@ VI.2 Relation to Prior Work
Watts (2002) [8] described stable nodes conceptually, and Gleeson & Cahalane (2007) [12] derived the analytical cascade condition; we operationalize the stable-node mechanism as the quantitative condition $k>1/\varphi$ and isolate its dynamical nature through controlled experiments that vary hub thresholds independently of topology. The hub vulnerability experiment (§III) provides a direct experimental decomposition of the stable-node mechanism from connectivity effects in the context of cascade dynamics.
@@
Gleeson (2008) [25] extended analytical cascade methods to correlated and modular networks in the static (unperturbed) case, providing the natural framework for extending our $z_{1}$ derivation to networks with degree-degree correlations—a direction we identify as important future work. Nishioka and Hasegawa (2022) [31] extended the Watts model with differential thresholds based on proximity to initiators, and Karimi and Holme (2021) [32] extended it to temporal networks. Our work complements these by studying how structural intervention (hub removal) modifies cascade thresholds, a distinct mechanism from seed selection or temporal dynamics. Lee et al. (2023) [34] studied threshold cascades on signed networks, demonstrating that heterogeneous tie types can shift phase boundaries—analogous to how hub removal shifts the cascade boundary in our setting.
Morone and Makse (2015) [16] and Braunstein et al. (2016) [17] optimize network dismantling for percolation fragmentation. Our results complement this literature by identifying the cascade side-effects of percolation-optimal interventions: dismantling strategies that efficiently fragment a network may simultaneously expand the cascade vulnerability window [27]. Geng et al. (2021) [30] proposed disassortative rewiring to enhance robustness of scale-free networks against localized attack, and Lv et al. (2022) [33] modeled SF network response to multi-node removal with dynamical cascading failures—both addressing the same hub-vulnerability mechanism from the defense perspective. The interdependent network literature [35, 38] has identified dual failure pathways in multi-layer settings; our contribution shows that compounding vulnerability arises even in single-layer networks when two stress models operate on the same topology, through a mechanism that recent hybrid spinodal analyses [36, 29] may help classify.
VI.3 Real-World Network Validation
To validate beyond synthetic BA networks, we test on four real-world networks from Stanford SNAP with diverse degree distributions and clustering coefficients, applying the same hub removal protocol ($\alpha=0.10$) and cascade sweep ($\varphi\in[0.05,0.40]$, 40 seeds per $\varphi$). We estimate $p_{c}$ using both the susceptibility-peak and $S=0.5$ methods.
Table 7: Real-world network validation of compounding vulnerability under hub removal ($\alpha=0.10$). $\Delta F_{p}$: percolation fragility change (susceptibility peak). $C_{\max}^{\mathrm{pre/post}}$: peak mean cascade size before/after hub removal across all $\varphi$ values tested.
Network NN $\kappa$ $C$ $\Delta F_{p}$ $C_{\max}^{\mathrm{pre}}$ $C_{\max}^{\mathrm{post}}$
p2p-Gnutella 6,299 17.7 0.011 $75\%$ $0.01\%$ $1.7\%$
Wiki-Vote 7,066 145.4 0.142 $594\%$ $0.01\%$ $6.0\%$
Bitcoin OTC 5,881 72.0 $170\%^{\ddagger}$ $0.06\%$ $0.8\%$
CA-HepPh 11,204 130.9 0.622 $58\%^{\ddagger}$ $0.04\%$ $2.9\%$
Bitcoin OTC has $\gamma<2$; the finite-$\kappa$ analytical framework is not strictly applicable. $S=0.5$ method used where susceptibility peak is unreliable. High clustering ($C=0.622$); susceptibility peak yields anomalous results.
[…]
Compounding vulnerability (both percolation and cascade worsening) is cleanly confirmed in p2p-Gnutella ($\kappa=17.7$, $C=0.011$; $\Delta F_{p}=+75\%$ by susceptibility peak) and Wiki-Vote ($\kappa=145.4$, $C=0.142$; $\Delta F_{p}=+594\%$). For Email-Enron ($C=0.497$) and CA-HepPh ($C=0.622$), the susceptibility-peak method gives anomalous percolation results ($\Delta F_{p}=-6.8\%$ and $80.4\%$ respectively), while the $S=0.5$ method yields $\Delta F_{p}=+31\%$ and $58\%$. We report $S=0.5$ values for these two networks (marked $\ddagger$} in Table 7). This estimator disagreement arises in high-clustering networks where the susceptibility landscape has multiple peaks; it does not affect the cascade amplification, which is robust across all estimators and all four networks. However, we note that full compounding vulnerability is confirmed with a single estimator only for the two low-clustering networks.
More broadly, our BA-derived $\kappa>{\sim}10$ threshold predicts compounding vulnerability in real-world networks: yeast protein-protein interaction networks ($\kappa\approx 17$ [23]), the World Wide Web ($\kappa\approx 55$ [24]), and citation networks ($\kappa>100$). Conversely, power grids ($\kappa\approx 2$) and regular-topology networks fall below the threshold.
@@ VI.5 Limitations
(i) We validate on synthetic networks with controlled degree heterogeneity and on multiple real-world networks from Stanford SNAP. The Bitcoin OTC Trust Network (N=5,881N=5{,}881, $\gamma\approx 1.9$) confirms the directional effect but, with $\gamma<2$ implying divergent $\langle k^{2}\rangle$, the finite-$\kappa$ analytical framework of §V is not strictly applicable to this network; the agreement is empirical rather than analytical. (ii) The bond percolation threshold $p_{c}$ is estimated using the susceptibility peak method (maximum of the mean finite-cluster size), the standard approach in computational percolation [7]. We verify robustness by comparing against the $S=0.5$ crossing and the Molloy–Reed analytical estimate; all three methods confirm $\Delta F_{p}\gg 0$ (Appendix A). (iii) We test sensitivity to threshold heterogeneity using Beta-distributed per-node thresholds centered on $\varphi$ (Appendix D); the compounding vulnerability effect persists under moderate and high heterogeneity, with amplification ranging from $349\%$ to $1{,}955\%$ at $\varphi=0.22$. (iv) The configuration-model approximation in §V.3 ignores BA disassortativity and the correlation structure induced by targeted removal. Appendix C validates the $z_{1}$ framework on configuration-model networks (where it is exact by construction), confirming that the approximation captures the dominant mechanism and that BA-specific correlations partially dampen rather than create the effect. (v) Static networks are studied; adaptive rewiring could mitigate or amplify the effect. (vi) The Watts cascade model is one of several cascade mechanisms; generalization to load-redistribution [18], cascade-based attacks [21], and DebtRank [14] models is an important direction. (vii) The $\kappa>{\sim}10$ threshold is derived for BA networks; the precise threshold depends on both the degree distribution shape and the cascade threshold $\varphi$, and should be understood as a BA-specific onset condition rather than a distribution-independent constant.
@@ VII Conclusion
We have demonstrated that in Barabási–Albert networks (N=2, [^0] N=2{,}000, $m=2$), hub removal can raise the bond percolation threshold from $p_{c}=0.174$ to $0.776$ while simultaneously expanding the Watts cascade window, enabling global cascades at threshold values that were previously subcritical. The phase transition at $\varphi=0.22$0.250.25—where mean cascade size jumps from $1\%$ to ${\sim}23\%$ of the network (1,0001{,}000 trials)—is the most striking consequence, and the vulnerability is latent: subcritical cascade behavior at these threshold values gives no warning of the system’s proximity to the phase boundary—a proximity that becomes critical only after hub removal shifts the system across it. In principle, a cascade-vulnerability diagnostic could detect this proximity, but in practice such diagnostics require knowledge of the post-intervention network, which is typically unavailable.
@@ Acknowledgements.
The author thanks the open-source Python community for NetworkX and NumPy. Simulation code, random seeds, and data are publicly available for independent reproduction [37].
@@ Appendix A Percolation Method Comparison
Table 8: Comparison of $p_{c}$ estimates across methods. “Susc. peak” is the susceptibility maximum (primary, Sec. II), “S=0.5S=0.5” is the giant component crossing, and “MR” is the Molloy–Reed analytical estimate.
Network $p_{c}$ (susc.) $p_{c}$ ($S=0.5$) $p_{c}$ (MR) $\Delta F_{p}$ (susc.)
BA before 0.174 0.337 0.092
BA after 0.776 0.903 0.240 $347\%$
ER before 0.354 0.251
WS before 0.499 0.323
@@ Appendix B $z_{1}$ Derivation Details
Appendix C Configuration-Model Validation
To validate the $z_{1}$ derivation (§V.3), which assumes configuration-model independence, we test on configuration-model networks constructed with the same degree sequence as BA(N=2, [^0] N=2{,}000, $m=2$). For each of 20 BA realizations, we generate a matched configuration-model network (removing self-loops and multi-edges), apply the same hub removal protocol, and compare cascade behavior.
Table 9: Configuration-model validation (N=2, [^0] N=2{,}000, $m=2$, $\varphi=0.22$, 20 networks $\times$ 20 seeds). The configuration model shows stronger cascades than BA, confirming that BA-specific degree correlations partially suppress the effect.
Network $\kappa$ $p_{c}^{\mathrm{pre}}$ $p_{c}^{\mathrm{post}}$ Post cascade
BA $12.0\pm 0.6$ 0.181 0.788 22.9%
Config model $11.4\pm 0.5$ 0.159 0.751 38.1%
The configuration-model networks exhibit stronger cascade amplification ($38.1\%$ vs $22.9\%$) and similar percolation degradation. This confirms two claims: (i) the $z_{1}$ framework accurately predicts cascade window expansion for configuration-model networks, where it is exact by construction; (ii) BA disassortativity partially dampens the cascade effect—high-degree hubs preferentially connect to low-degree stubs, creating extra “firewall” correlations absent in uncorrelated networks. The analytical upper-bound character of Eq. 3 is thus validated: the true BA cascade size lies below the configuration-model prediction.
Appendix D Heterogeneous Threshold Sensitivity
To test whether compounding vulnerability depends on the assumption of uniform thresholds, we replace the fixed $\varphi=0.22$ with per-node thresholds drawn from three distributions, each centered on $\varphi=0.22$: (i) Uniform($0.12,0.32$), (ii) Beta with low heterogeneity ($\alpha=5,\beta=17.7$, $\sigma\approx 0.08$), and (iii) Beta with high heterogeneity ($\alpha=2,\beta=7.1$, $\sigma\approx 0.13$). Each condition uses 600 trials (15 networks $\times$ 40 seeds).
Table 10: Heterogeneous threshold sensitivity ($\varphi=0.22$, 600 trials each). “Amp.” is the amplification at $\varphi=0.22$; the maximum across all $\varphi$ (peaking at $\varphi\approx 0.24$) is larger in all cases.
Distribution Pre (%) Post (%) Amp.
Uniform (homogeneous) 1.0 21.6 $1{,}955\%$
Beta (low het.) 1.3 25.1 $1{,}830\%$
Beta (high het.) 4.0 18.0 $349\%$
Under all three distributions, hub removal produces massive cascade amplification at the transition ($349\%$ to $1{,}955\%$ at $\varphi=0.22$; amplification peaks at $\varphi\approx 0.24$ in all cases). High heterogeneity reduces the effect because some nodes have thresholds well below 0.22, making them vulnerable even when hubs are present—partially pre-empting the firewall destruction mechanism. Critically, the qualitative effect (subcritical $\to$ supercritical transition) persists: compounding vulnerability is not an artifact of uniform thresholds.
Appendix E Simulation Code and Formal Proofs
@@
A Lean 4 formalization of key analytical claims from §V is provided in proofs/ in the same repository, including the stable-node condition (Definition 1), the cascade onset bound (Observation 1), and the monotonicity of $p_{c}$ under hub removal (Proposition 1). These machine-verified proofs establish the mathematical correctness of the cascade-blocking inequalities independently of the simulation results.
Key scripts: simulate_production.py (finite-size scaling), phi_sweep_large.py ($\varphi$-sweep), hub_vulnerability_test.pyIII experiment), heterogeneous_threshold_test.py (Appendix D).